### v1

@@ Appendix B Laboratory instrument characterization
Before deployment at the telescope, the instrument was extensively tested in laboratory environment. Performed measurements and analysis include optical quality and alignment, instrument sensitivity, and stability, among others. A dedicated experimental setup was built to mimic the instrument operational conditions. The system includes a mechanical structure to tilt the instrument, equivalent to 60-90 deg of elevation at telescope; a wire scanner to generate small signals over a cold background load;and a broad band photonic continuous wave signal generator mounted on a X-Y stage allowing for beam characterization of the system, as described in Reyes_2023.
@@
Of particular interest is the wire scanner device, presented in section 9. It consist of two small wires that moves against the sky background. Position of the wires is accurately record using Precision Time Protocol (PTP) over Ethernet and synchronized with data acquisition. The wire size (0.5mm) and material(standard electronic wire, plastic coated) were chosen to generate a small 10 K signal against the cold background.
@@
The wire scanner was originally developed for fast characterization of the instrument performance in laboratory enviroment. For this a cold load is placed behind the wires mimicking the sky background. The load wass implemented as a large copper bath covering the complete focal plane (80x80 cm2). The bucket lower side is externally coated with Stycast and silicon carbide grains to act as a cold load. The bath is thermally insulated by placing it inside a larger 4 cm wall thickness styrofoam box. The effective temperature of this cold load was measured using a KID detector together with a rotating grid. The grid switches between two standard temperatures, ambient load at 300 K and liquid nitrogen load at 80 K. Measurements indicate an effective temperature of the load of 120 K for the low frequency band 160 K for the high frequency band, elevated from 77K due to the styrofoam losses. However, these values are similar to the expected sky temperature at the deployment site.
@@
Appendix C Magnetic shielding
@@
The low frequency array presented a prominent 0.3 Hz signal over most of the array, but specially strong at chip edges. The amplitud of the signal was as large as 0.5 K. This signal is most likely magnetically coupled to the detectors.
@@
This behavior presented a challenge for operations, but as induced signals are slow and correlated over the array they can be partially filtered out in post-processing. A more serious problem is the hysteretic response of some detectors. This is seen as a sudden jump of level of the timelines of some detectors, usually happening when magnetic flux exceeds a threshold as described in Stan_2004. This high frequency content signal cannot be filtered out and the affected detectors cannot be used for astronomical observations. This topic was investigated experimentally in laboratory environment by mounting an external small coil on the cryostat ( see Fig. 17) to artificially generate magnetic field as expected in regular operations. The main conclusion of the study is that the HFA performs well when facing magnetic fields up to 15 T (maximal change during an Azimuth scan due to earth magnetic field). For the same field intensity the LFA presents a hysteric behavior. This noticeable difference is due to the increased pixel size at low frequency, giving a larger focusing of flux into the sensitive part of the KIDs. Other possible reasons for this difference are the higher kinetic inductance of the LFA KIDs due to the thinner Aluminum film, and the slightly worst shielding of the LFA by the Aluminum cup.
@@
Appendix D The readout noise contribution
The digitization noise caused by the finite number of bits in the ADC and DAC can be expressed by the power spectral density (PSD) of the read tone normalized by the off resonance tone power as:
where GSamples/s is the sampling rate of the ADC; is the noise degradation due to the number of tones (n) used by the read-out; is the crest factor, the ratio between the peak voltage and the root mean square voltage of the waveform. We use a random initialization of tone phases achieving a typical crest factor of 14 dB. The factor 3 dB is because of the the double sideband down converter, which increases the noise floor by a factor of 2. The effective number of bits (ENOB) is a measure of the real performance of the digitization hardware and is lower than the real number of bits. The AMKID read-out uses a 14 bit DAC and a 10 bit ADC, being the total PSD of the system dominated by the ADC noise, characterized by an ENOB of 7.6dB.
For 1000 tones the read-out achieves a of -93.8 dBc/Hz, assuming flat power across the band.
The system noise analysis shows that our read-out noise is dominated by the above described digitization noise and not by the analog section of the read-out, being the last more than 2 order of magnitude smaller that the digitization noise (-125.2 dBc/Hz).
The digitized signal is integrated in a circular buffer and later Fourier transform as explained in section 8. The output signal from the digital section is the gain coefficient associate to each readout tone. We can express the digitization noise as a standard deviation of the measured gain coefficient in a one second measurement as.
@@ Appendix D The readout noise contribution
Where K is the conversion factor from phase shift () to temperature scale as discussed in section 9; fsamp the sampling frequency; and R is the radius of the KID circle, which is related to the KID depth (min(S) as:
@@
In our system we obtain a typical separation (ratio between and ) of 2.8 dB, meaning that the ADC contribution is on the same order of magnitude than the instrument NET. Giving the cuadratic adding of NET this means that the ADC noise accounts for a 15% of the measured NET, indicating a clear path to improve the system performance by using a better read-out with a larger ENOB.
Appendix E NET modelling
@@
The number of photons per unit of frequency that reach a detector during a second is calculated as:
From here we calculate the Noise Equivalent power (NEP) of a single detector as :
Expression that include the Poisson noise together with the wave and recombination noise contributions, see Flanigan_2016 and Ferrari_2018a. is the efficiency of the quasiparticle generation and is the superconductor bandgap frequency. The factor is used to express the NEP in units of , i.e. as the S/N obtained in a 0.5 s integration time.
@@
Where is the Boltzmann constant and the filter bandpass. The factor is used here to express the NET in units of , i.e. as S/N after a 1 second integration time. In our case the factor 2 in the denominator cancels with the , but it is explicit in the equation to facilitate extension of this procedure for dual polarization instruments.

### v2

@@ Appendix A Laboratory instrument characterization
Before deployment at the telescope, the instrument was extensively tested in laboratory environment. Performed measurements and analysis include optical quality and alignment, instrument sensitivity, and stability, among others. A dedicated experimental setup was built to mimic the instrument operational conditions. The system includes a mechanical structure to tilt the instrument, equivalent to 60-90 deg of elevation at telescope; a wire scanner to generate small signals over a cold background load; and a broad band photonic continuous wave signal generator mounted on a X-Y stage allowing for beam characterization of the system, as described in Reyes et al. (2023).
@@
Of particular interest is the wire scanner device, presented in Sect. 10. It consist of two small wires that moves against the sky background. Position of the wires is accurately record using Precision Time Protocol (PTP) over Ethernet and synchronized with data acquisition. The wire size (0.5mm) and material(standard electronic wire, plastic coated) were chosen to generate a small 10 K signal against the cold background.
@@
The wire scanner was originally developed for fast characterization of the instrument performance in laboratory environment. For this a cold load is placed behind the wires mimicking the sky background. The load was implemented as a large copper bath covering the complete focal plane (80x80 cm2). The bucket lower side is externally coated with Stycast conductive epoxy and silicon carbide grains to act as a cold load. The bath is thermally insulated by placing it inside a larger 4 cm wall thickness Styrofoam box. The effective temperature of this cold load was measured using a KID detector together with a rotating grid. The grid switches between two standard temperatures, ambient load at 300 K and liquid nitrogen load at 80 K. Measurements indicate an effective temperature of the load of 120 K for the low frequency band 160 K for the high frequency band, elevated from 77K due to the Styrofoam losses. However, these values are similar to the expected sky temperature at the deployment site.
@@
Appendix B Magnetic shielding
@@
The low frequency array presented a prominent 0.3 Hz signal over most of the array, but specially strong at chip edges. The amplitud of the signal was as large as 0.5 K, as seen in Fig 18. This signal is most likely magnetically coupled to the detectors.
@@
This behavior presented a challenge for operations, but as induced signals are slow and correlated over the array they can be partially filtered out in post-processing. A more serious problem is the hysteric response of some detectors. This is seen as a sudden jump of level of the timelines of some detectors, usually happening when magnetic flux exceeds a threshold as described in Stan et al. (2004). This high frequency content signal cannot be filtered out and the affected detectors cannot be used for astronomical observations. This topic was investigated experimentally in laboratory environment by mounting an external small coil on the cryostat ( see Fig. 17) to artificially generate magnetic field as expected in regular operations. The main conclusion of the study is that the HFA performs well when facing magnetic fields up to 15 T (maximal change during an Azimuth scan due to earth magnetic field). For the same field intensity the LFA presents a hysteric behavior. This noticeable difference is due to the increased pixel size at low frequency, giving a larger focusing of flux into the sensitive part of the KIDs. Other possible reasons for this difference are the higher kinetic inductance of the LFA KIDs due to the thinner aluminum film, and the slightly worst shielding of the LFA by the aluminum cup.
@@
Appendix C The readout noise contribution
The digitization noise caused by the finite number of bits in the ADC and DAC can be expressed by the PSD of the read tone normalized by the off-resonance tone power as
where GSamples/s is the sampling rate of the ADC; is the noise degradation due to the number of tones (n) used by the readout; is the crest factor, the ratio between the peak voltage and the root mean square voltage of the waveform. We use a random initialization of tone phases achieving a typical crest factor of 14 dB. The factor 3 dB is because of the the double sideband down converter, which increases the noise floor by a factor of 2. The effective number of bits (ENOB) is a measure of the real performance of the digitization hardware and is lower than the real number of bits. The AMKID readout uses a 14-bit DAC and a 10-bit ADC, being the total PSD of the system dominated by the ADC noise, characterized by an ENOB of 7.6dB.
For 1000 tones the readout achieves a of -93.8 dBc/Hz, assuming flat power across the band.
The system noise analysis shows that our readout noise is dominated by the above described digitization noise and not by the analog section of the readout, being the last more than 2 order of magnitude smaller that the digitization noise (-125.2 dBc/Hz).
The digitized signal is integrated in a circular buffer and later Fourier transform as explained in Sect. 9. The output signal from the digital section is the gain coefficient associate to each readout tone. We can express the digitization noise as a standard deviation of the measured gain coefficient in a one second measurement as
@@ Appendix C The readout noise contribution
where K is the conversion factor from phase shift () to temperature scale as discussed in Sect. 10; fsamp the sampling frequency; and R is the radius of the KID circle, which is related to the KID depth (min(S) as
@@
In our system we obtain a typical separation (ratio between and ) of 2.8 dB, meaning that the ADC contribution is on the same order of magnitude than the instrument NET. Giving the quadratic adding of NET this means that the ADC noise accounts for a 15% of the measured NET, indicating a clear path to improve the system performance by using a better readout with a larger ENOB.
Appendix D NET modeling
@@
The number of photons per unit of frequency that reach a detector during a second is calculated as
From here we calculate the Noise Equivalent power (NEP) of a single detector as
Expression that include the Poisson noise together with the wave and recombination noise contributions, see Flanigan et al. (2016) and Ferrari et al. (2018b). is the efficiency of the quasiparticle generation and is the superconductor bandgap frequency. The factor is used to express the NEP in units of , i.e., as the S/N obtained in a 0.5 s integration time.
@@
where is the Boltzmann constant and the filter bandpass. The factor is used here to express the NET in units of , i.e., as S/N after a 1 second integration time. In our case the factor 2 in the denominator cancels with the , but it is explicit in the equation to facilitate extension of this procedure for dual polarization instruments.